# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v4

@@ 5 Output
The zero terms are from perfect indistinguishability (Lemmas P.7 and P.9).
@@
    Privacy: Response shares $\mathbf{z}_{i}$ enjoy statistical privacy (Theorem 4.4) even with a single honest signer; commitment and r0-check values hidden by pairwise masks if $\geq 2$ honest signers
@@
Privacy: By Lemma P.9, honest signers’ responses have high conditional min-entropy (statistical, via Shamir nonce DKG). The $\mathbf{V}_{j}$ values are computationally hidden by pairwise masks. The 2PC reveals only the result bit. Combining: adversary view simulatable from public values + corrupted shares + result bit.
Liveness: The protocol has explicit timeout. If conditions met, 2PC completes in bounded time (synchronous between CPs). If conditions not met, protocol aborts (no liveness violation, just unsuccessful signing).
@@
Proof P.16.
Signer experience: After receiving challenge, signer computes response (local) and sends once. No further interaction required.
Server interaction: CPs execute 2PC (constant rounds with GC precomputation). With precomputation, effectively 1 round (OT + evaluate).
Retry handling: Each retry requires fresh pre-signing set and response. Signer daemon handles automatically; user sees only final result.
@@
Profile Rounds/Attempt Expected Attempts Total Rounds
@@
Table 16: Round complexity comparison across profiles. P3 denotes the synchronous 2PC variant (5 rounds) prior to the semi-async optimization; P3+ is the final semi-async protocol presented in Section 6.5.
@@
Let $S\subseteq[N]$ be a signing set with $|S|\geq 2$. For any domain separator $\mathsf{dom}$ and PRF $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{d}$, define:
$\mathbf{m}_{i}=\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{dom})$
Then $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}\in R_{q}^{d}$.
@@
For any pair $(i,j)\in S\times S$ with $i<j$, the term $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})$ appears:
@@
$\sum_{i\in S}\mathbf{m}_{i}$ $=\sum_{i\in S}\left(\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{dom})\right)$
$=\sum_{\begin{subarray}{c}(i,j)\in S\times S\\ i<j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})-\sum_{\begin{subarray}{c}(i,j)\in S\times S\\ i<j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})$
@@
Let $S$ be a signing set with $|S|\geq T$, let $C\subset S$ with $|C|\leq T-1$ be the set of corrupted parties, and let $h\in S\setminus C$ be an honest party. In the Shamir nonce DKG protocol, the adversary’s view of $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ has conditional min-entropy:
$H_{\infty}\!\left(\mathbf{z}_{h}\mid\{\mathbf{f}_{j}\}_{j\in C},\;\{\mathbf{f}_{h}(j)\}_{j\in C},\;c\right)\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
No computational assumptions are needed. Note that the conditioning does not include individual $\mathbf{W}_{i}=\lambda_{i}\mathbf{A}\mathbf{y}_{i}+\mathbf{m}_{i}^{(w)}$ values, which are sent only to the coordinator (not broadcast) in the protocol. Including the coordinator’s broadcast $\mathbf{w}_{1}=\mathsf{HighBits}(\sum_{i}\mathbf{W}_{i},2\gamma_{2})$ does not reduce the bound, since $\mathbf{w}_{1}$ is already implicit in any valid signature. For ML-DSA-65 with $|S|\leq 17$, this exceeds 24,00024{,}000 bits—over $5\times$ the secret key entropy $H(\mathbf{s}_{1,h})\leq 4058$ bits.
@@
The proof proceeds in four steps, following a null space argument on the honest party’s polynomial.
The adversary’s knowledge about the honest party’s polynomial $\mathbf{f}_{h}$ consists of the evaluations $\{\mathbf{f}_{h}(j)\}_{j\in C}$ received during the DKG. Since $|C|\leq T-1$ and $\mathbf{f}_{h}$ has degree T1T-1 (hence $T$ coefficients), the adversary has at most T1T-1 linear equations in $T$ unknowns.
Step 1: Null space. The evaluation map $\mathbf{f}_{h}\mapsto(\mathbf{f}_{h}(j_{1}),\ldots,\mathbf{f}_{h}(j_{|C|}))$ has a null space containing the polynomial $\mathbf{n}(x)=\prod_{j\in C}(x-x_{j})$, which has degree $|C|\leq T-1$ and leading coefficient 1. The set of polynomials consistent with the adversary’s observations is $\mathbf{f}_{h}(x)=\mathbf{g}(x)+\mathbf{t}\cdot\mathbf{n}(x)$, where $\mathbf{g}$ is any fixed interpolant through the known points and $\mathbf{t}\in R_{q}^{\ell}$ is the free parameter.
Step 2: Posterior constraint on $\mathbf{t}$. The coefficient $\mathbf{a}_{h,T-1}$ is sampled uniformly from $R_{q}^{\ell}$ a priori. However, the constant term $\hat{\mathbf{y}}_{h}=\mathbf{f}_{h}(0)=\mathbf{g}(0)+\mathbf{t}\cdot\mathbf{n}(0)$ must satisfy $\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$. Since $\mathbf{n}(0)\neq 0\pmod{q}$, this constrains each coefficient of $\mathbf{t}$ to at most $2\lfloor\gamma_{1}/|S|\rfloor+1$ values. The posterior distribution of $\mathbf{t}$ is not $\mathbb{Z}_{q}$-uniform, but retains per-coefficient support of this size.
@@
Step 4: Response. The response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ shifts again by a fixed quantity, yielding $H_{\infty}(\mathbf{z}_{h}\mid\mathsf{View}_{C})\geq n\ell\cdot\log_{2}(2\lfloor\gamma_{1}/|S|\rfloor+1)$.
@@
Note: This lemma applies to the commitment and r0-check values, which still use pairwise-canceling masks. The response $\mathbf{z}_{j}=\mathbf{y}_{j}+c\cdot\mathbf{s}_{1,j}$ achieves statistical privacy via the Shamir nonce DKG (Theorem Q.3) and does not require masks.
@@
$\mathbf{m}_{j}^{(s2)}$ $=\underbrace{\sum_{k\in C\cap S:k>j}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})-\sum_{k\in C\cap S:k<j}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{dom})}_{\text{computable by }\mathcal{A}}$
$\quad+\underbrace{\sum_{k\in(S\setminus C)\setminus\{j\}:k>j}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})-\sum_{k\in(S\setminus C)\setminus\{j\}:k<j}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{dom})}_{\text{unknown to }\mathcal{A}}$
Since $|S\setminus C|\geq 2$, there exists at least one honest party $k^{\prime}\in(S\setminus C)\setminus\{j\}$ whose PRF seed $\mathsf{seed}_{j,k^{\prime}}$ is unknown to $\mathcal{A}$. By PRF security, $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{dom})$ is computationally indistinguishable from uniform in $R_{q}^{k}$. Since the mask contains at least one such uniform-looking term, and uniform plus fixed equals uniform, $\mathbf{V}_{j}\stackrel{{\scriptstyle c}}{{\approx}}\mathsf{Uniform}(R_{q}^{k})$. The distinguishing advantage is bounded by $|S\setminus C|\cdot\epsilon_{\mathsf{PRF}}$.
@@
$\mathsf{Var}(X_{i})$ $=\mathbb{E}[X_{i}^{2}]=\frac{1}{2B+1}\sum_{k=-B}^{B}k^{2}=\frac{1}{2B+1}\cdot\frac{2B(B+1)(2B+1)}{6}=\frac{B(B+1)}{3}$
@@
For $\epsilon\geq 0$:
$R_{\alpha}^{\epsilon}(P\|Q)=\min_{\begin{subarray}{c}P^{\prime}:\Delta(P^{\prime},P)\leq\epsilon\\ Q^{\prime}:\Delta(Q^{\prime},Q)\leq\epsilon\end{subarray}}R_{\alpha}(P^{\prime}\|Q^{\prime})$
where $\Delta$ denotes statistical distance.
@@
Let $\Pi_{\mathsf{IH}}$ be threshold ML-DSA with Irwin-Hall nonces (sum of $|S|$ uniforms), and $\Pi_{\mathsf{U}}$ be the hypothetical scheme with uniform nonces. Then for any PPT adversary $\mathcal{A}$:
$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}(\mathcal{A})}+q_{S}\cdot\epsilon$
where $q_{S}$ is the number of signing queries, $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the full-vector Rényi divergence, and the per-coordinate $R_{2}^{\epsilon,\mathsf{coord}},\epsilon$ are computed below.
Proof Q.14.
@@
In $\Pi_{\mathsf{IH}}$ (real): $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ where each $\hat{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$ is party $i$’s constant-term contribution in the nonce DKG.
@@
For $n=|S|$ uniforms on range $N=2\lfloor\gamma_{1}/|S|\rfloor$ with shift $c=\|(c\mathbf{s}_{1})_{j}\|_{\infty}\leq\beta$, specializing to $\alpha=2$ (where the factor $\alpha(\alpha-1)/2=1$):
$R_{2}^{\epsilon}\leq 1+\frac{n^{2}c^{2}}{N^{2}}+O(1/n!)$

### v5

@@ 5 Output
The hybrid $\mathsf{H}_{2.5,\mathsf{ext}}$ (between $\mathsf{H}_{2}$ and $\mathsf{H}_{3}$) is the extraction hybrid: when $\mathsf{CP}_{1}$ is corrupted, the simulator uses LP07 cut-and-choose extractability to extract $\mathbf{S}_{1}$ from the adversarially-submitted garbled circuit; this succeeds except with probability $\epsilon_{\mathsf{ext}}\leq 2^{-B/2}$ (see the extraction error accounting below). The three subsequent zero terms arise from perfect indistinguishability: pre-signing simulation (Lemma P.8), response simulation (Lemma P.10), and the transition to the ideal execution.
Hint output and combiner role. The ideal functionality $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ (Definition P.3) outputs both the 1-bit result and the hint $\mathbf{h}=\mathsf{MakeHint}(\mathbf{w}^{\prime},\mathbf{w},2\gamma_{2})$ to the combiner. In both simulator cases, $\mathcal{S}^{\mathsf{P3+}}$ receives $(\mathsf{result},\mathbf{h})$ from $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ and passes $\mathbf{h}$ to the adversary. The hint $\mathbf{h}$ appears in the public signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ and creates no leakage beyond what $\sigma$ already reveals; the BHR12 simulator constructs a garbled circuit consistent with output $(\mathsf{result},\mathbf{h})$ without exposing $c\mathbf{s}_{2}$.
Combiner and CP roles. In P3+, $\mathsf{CP}_{1}$ (the garbler) also acts as the combiner: it aggregates signer responses $(\mathbf{z}_{j},\mathbf{V}_{j})$, computes the split $(S_{1},S_{2})$ for the 2PC, and executes the garbled-circuit protocol with $\mathsf{CP}_{2}$. Combiner corruption is subsumed by the $\mathsf{CP}_{1}$ is corrupted” case (Simulator P.6): the simulator uses LP07 extraction to recover $\mathbf{S}_{1}$ from the adversary’s garbled circuit and forwards honest parties’ contributions via the ideal-world channel to $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$.
Extraction error accounting. When $\mathsf{CP}_{1}$ (garbler) is corrupted, $\mathcal{S}^{\mathsf{P3+}}$ uses LP07 extractability to recover $\mathbf{S}_{1}$ from the adversarially-submitted garbled circuit. The cut-and-choose protocol (generate $B$ circuits; open $\lfloor B/2\rfloor$ for verification; use the rest) ensures a cheating garbler is caught with probability $1-2^{-\lfloor B/2\rfloor}\geq 1-2^{-B/2}$. For unopened circuits, LP07 extraction succeeds with probability $1$ (requires only correct evaluation, guaranteed by cut-and-choose). Thus $\epsilon_{\mathsf{ext}}\leq 2^{-B/2}$; setting $B=2\kappa$ gives $\epsilon_{\mathsf{ext}}\leq 2^{-\kappa}$. The full bound $\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{GC}}+2^{-B/2}<2^{-\kappa+1}$ holds for standard $\kappa$-bit-secure OT and GC.
@@
    Privacy: Nonce shares $\mathbf{y}_{i}$ enjoy statistical privacy (Theorem 4.7); key privacy given $\sigma$ is conjectured to hold computationally under Module-LWE (open problem); commitment and r0-check values hidden by pairwise masks if $\geq 2$ honest signers
@@
Privacy: By Lemma P.10, the simulator computes honest signers’ responses from simulated DKG values (perfect indistinguishability). Nonce shares have high conditional min-entropy (Theorem 4.7, statistical); key privacy given $\sigma$ is conjectured to hold computationally under Module-LWE (formal reduction is an open problem, unresolved even in the single-signer setting). The $\mathbf{V}_{j}$ values are computationally hidden by pairwise masks. The 2PC reveals the result bit and hint $\mathbf{h}$ (both appear in the public signature); $\mathbf{h}$ does not create additional privacy leakage beyond what $\sigma$ already reveals. Combining: adversary view simulatable from public values + corrupted shares + $(\mathsf{result},\mathbf{h})$.
Liveness: The protocol has an explicit timeout. If conditions are met, the 2PC completes in bounded time (synchronous between CPs). If not, the protocol aborts (no liveness violation; just unsuccessful signing).
@@
Proof P.17.
Signer experience: After receiving the challenge, the signer computes a local response and sends it once. No further interaction is required.
Server interaction: CPs execute 2PC in constant rounds (with GC precomputation), effectively 1 round (OT + evaluate).
Retry handling: Each retry requires a fresh pre-signing set and response; the signer daemon handles this automatically.
@@
Profile Rds/Attempt Exp. Attempts Total Rds
@@
Table 16: Round complexity across profiles. P3: synchronous 2PC variant (5 rounds) before the semi-async optimization; P3+ is the final protocol (Section 6.5).
@@
Let $S\subseteq[N]$ be a signing set with $|S|\geq 2$. For any PRF input $\mathsf{ctx}\in\{0,1\}^{*}$ (typically $\mathsf{ctx}=\mathsf{session\_id}\|\mathsf{dom}$, where $\mathsf{session\_id}$ is a fresh per-session identifier and $\mathsf{dom}$ is a fixed domain label) and PRF $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{k}$ (where $k=6$ is the module rank; not the dropping-bits parameter $d=13$), define:
$\mathbf{m}_{i}=\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{ctx})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{ctx})$
Then $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}\in R_{q}^{k}$.
@@
For any pair $(i,j)\in S\times S$ with $i<j$, the term $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{ctx})$ appears:
@@
$\sum_{i\in S}\mathbf{m}_{i}$ $=\sum_{i\in S}\left(\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{ctx})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{ctx})\right)$
$=\sum_{\begin{subarray}{c}(i,j)\in S\times S\\ i<j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{ctx})-\sum_{\begin{subarray}{c}(i,j)\in S\times S\\ i<j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{ctx})$
@@
Let $S$ be a signing set with $|S|\geq T$, let $C\subset S$ with $|C|\leq T-1$ be the set of corrupted parties, and let $h\in S\setminus C$ be an honest party. In the Shamir nonce DKG protocol (Algorithm 2), each party samples its constant-term contribution $\hat{\mathbf{y}}_{h}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$ and constructs a degree-(T1)(T-1) polynomial with uniform higher-degree coefficients. The honest party’s nonce share $\mathbf{y}_{h}=\sum_{i\in S}\mathbf{f}_{i}(h)$ has conditional min-entropy:
$H_{\infty}\!\left(\mathbf{y}_{h}\mid\{\mathbf{f}_{j}\}_{j\in C},\;\{\mathbf{f}_{h}(j)\}_{j\in C}\right)\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
where the inequality holds directly from the sampling procedure (no rejection conditioning needed): the bounded sampling range for each party’s constant term directly ensures that the honest party’s polynomial $\mathbf{f}_{h}$ has $\mathbf{f}_{h}(0)=\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$ by construction (Algorithm 2, Line 2). No computational assumptions are needed for the nonce privacy; it is purely statistical. For ML-DSA-65 with $|S|\leq 17$, this exceeds 20,00020{,}000 bits—over $5\times$ the short secret key entropy $H(\mathbf{s}_{1})\leq 4058$ bits. (Shamir shares $\mathbf{s}_{1,h}$ for $h\neq 0$ are $\mathbb{Z}_{q}$-uniform; the entropy comparison is against the short secret $\mathbf{s}_{1}=\mathbf{f}(0)$.)
Note on aggregated nonce $\mathbf{y}=\sum_{i}\hat{\mathbf{y}}_{i}$. The aggregated nonce $\mathbf{y}$ (sum of all parties’ constant terms) follows an Irwin-Hall distribution with tails extending beyond $\gamma_{1},\gamma_{1}]$. However, the aggregated nonce is never required to lie strictly within this range for the min-entropy bound to hold: the bound applies to individual shares $\mathbf{y}_{h}$, not the aggregate. The z-bound check $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ implicitly rejects sessions where the aggregate $\|\mathbf{y}\|_{\infty}$ is too large (via $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$), but this rejection event is already accounted for in the Irwin-Hall tail probability $\epsilon<10^{-30}$ (Lemma Q.22); it does not affect the per-share conditional min-entropy.
Key privacy given $\sigma$ (computational, informal). Even given the public signature $\sigma$ (from which $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ is computable when $|S\setminus C|=1$), extracting $\mathbf{s}_{1,h}$ requires solving a bounded-distance decoding problem with sparse multiplier $c$. While not a standard Module-LWE instance, this is at least as hard as key recovery from single-signer ML-DSA signatures; see the discussion in Theorem 4.7 (Section 4).
@@
The proof proceeds in three steps, following a null space argument on the honest party’s polynomial.
Reduction to the worst case |C|=T1|C|=T-1. The stated min-entropy lower bound is monotone non-decreasing as $|C|$ decreases: fewer corrupted evaluations give a higher-dimensional null space (dimension T|C|>1T-|C|>1 when |C|<T1|C|<T-1), meaning the adversary is more uncertain about $\mathbf{y}_{h}$ (higher conditional min-entropy). Hence it suffices to prove the bound for the hardest case |C|=T1|C|=T-1 (minimum uncertainty for the adversary); the case |C|<T1|C|<T-1 gives at least as much privacy and follows a fortiori. We henceforth assume |C|=T1|C|=T-1.
The adversary’s knowledge about the honest party’s polynomial $\mathbf{f}_{h}$ consists of the evaluations $\{\mathbf{f}_{h}(j)\}_{j\in C}$ received during the DKG. With |C|=T1|C|=T-1 and $\mathbf{f}_{h}$ having degree T1T-1 (hence $T$ coefficients), the adversary has exactly T1T-1 linear equations in $T$ unknowns: the null space of the evaluation map has dimension exactly $1$.
Step 1: Null space. The evaluation map $\mathbf{f}_{h}\mapsto(\mathbf{f}_{h}(j_{1}),\ldots,\mathbf{f}_{h}(j_{T-1}))$ (with |C|=T1|C|=T-1) has a one-dimensional null space, spanned by the monic polynomial $\mathbf{n}(x)=\prod_{j\in C}(x-x_{j})$ of degree T1T-1 and leading coefficient $1$. The set of polynomials consistent with the adversary’s observations is $\mathbf{f}_{h}(x)=\mathbf{g}(x)+\mathbf{t}\cdot\mathbf{n}(x)$, where $\mathbf{g}$ is any fixed interpolant through the known points and $\mathbf{t}\in R_{q}^{\ell}$ is the unique free parameter.
Step 2: Posterior constraint on $\mathbf{t}$. In the nonce DKG, each honest party $h$ samples $\hat{\mathbf{y}}_{h}$ directly from $\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$ and all T1T-1 higher-degree coefficients $\mathbf{a}_{h,1},\ldots,\mathbf{a}_{h,T-1}$ independently and uniformly from $R_{q}^{\ell}$. In the null-space parameterization, the free parameter is $\mathbf{t}=\mathbf{a}_{h,T-1}-\mathbf{g}_{T-1}$, where $\mathbf{g}_{T-1}$ is the leading coefficient of the fixed interpolant $\mathbf{g}$. Because $\mathbf{a}_{h,T-1}$ is $\mathbb{Z}_{q}$-uniform a priori, $\mathbf{t}$ is $\mathbb{Z}_{q}$-uniform per coefficient. The constant term $\hat{\mathbf{y}}_{h}=\mathbf{f}_{h}(0)=\mathbf{g}(0)+\mathbf{t}\cdot\mathbf{n}(0)$ must satisfy $\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$. Since $\mathbf{n}(0)=\prod_{j\in C}(-j)\neq 0\pmod{q}$ (each $j\in C$ satisfies $1\leq j\leq N\leq q-1$, so $j\not\equiv 0\pmod{q}$, and the product of nonzero field elements is nonzero), multiplication by $\mathbf{n}(0)$ is a bijection on $\mathbb{Z}_{q}$. Since the Shamir evaluation map acts independently on each of the $n\ell=1280$ coefficient slots (each coefficient of $\mathbf{f}_{h}$ follows a separate scalar Shamir sharing over $\mathbb{Z}_{q}$), this bijection applies coordinate-by-coordinate across all $n\ell$ slots; the constraint selects exactly $2\lfloor\gamma_{1}/|S|\rfloor+1$ values per coefficient of $\mathbf{t}$. Because the prior is $\mathbb{Z}_{q}$-uniform, conditioning on this subset yields a uniform posterior over the valid values.
@@
Key privacy (computational). Given the public signature $\sigma$, the adversary can compute $\mathbf{z}_{h}$ (when $|S\setminus C|=1$). The equation $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ relates two bounded unknowns. The high min-entropy of $\mathbf{y}_{h}$ ensures this is a non-trivial bounded-distance decoding instance: extracting $\mathbf{s}_{1,h}$ is computationally hard under the Module-LWE assumption.
@@
Note: This lemma applies to the commitment and r0-check values, which still use pairwise-canceling masks. The response $\mathbf{z}_{j}=\mathbf{y}_{j}+c\cdot\mathbf{s}_{1,j}$ does not require masks: the underlying nonce share $\mathbf{y}_{j}$ has statistical privacy via the Shamir nonce DKG (Theorem Q.3), enabling perfect simulation of $\mathbf{z}_{j}$ from simulated DKG values.
@@
$\mathbf{m}_{j}^{(s2)}$ $=\underbrace{\sum_{\begin{subarray}{c}k\in C\cap S\\ k>j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{ctx})-\sum_{\begin{subarray}{c}k\in C\cap S\\ k<j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{ctx})}_{\text{known to }\mathcal{A}}$
$\quad+\underbrace{\sum_{\begin{subarray}{c}k\in(S\setminus C)\setminus\{j\}\\ k>j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{ctx})-\sum_{\begin{subarray}{c}k\in(S\setminus C)\setminus\{j\}\\ k<j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{ctx})}_{\text{unknown to }\mathcal{A}}$
Since $|S\setminus C|\geq 2$, there exists at least one honest party $k^{\prime}\in(S\setminus C)\setminus\{j\}$ whose PRF seed $\mathsf{seed}_{j,k^{\prime}}$ is unknown to $\mathcal{A}$. By PRF security, $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{ctx})$ is computationally indistinguishable from uniform in $R_{q}^{k}$. Since the mask contains at least one such uniform-looking term, and uniform plus fixed equals uniform, $\mathbf{V}_{j}\stackrel{{\scriptstyle c}}{{\approx}}\mathsf{Uniform}(R_{q}^{k})$. One unknown PRF term suffices; the distinguishing advantage is bounded by $\epsilon_{\mathsf{PRF}}$ (or at most $(|S\setminus C|-1)\cdot\epsilon_{\mathsf{PRF}}$ via a union bound over honest-honest seed pairs).
@@
$\mathsf{Var}(X_{i})$ $=\mathbb{E}[X_{i}^{2}]=\frac{1}{2B+1}\sum_{k=-B}^{B}k^{2}$
$=\frac{1}{2B+1}\cdot\frac{2B(B+1)(2B+1)}{6}=\frac{B(B+1)}{3}$
@@
We use the exponential (non-log) form throughout, following [BLL+15, Raccoon2024]. In this convention $R_{2}(P\|Q)=\sum_{x}P(x)^{2}/Q(x)$, the product rule $R_{2}(P_{1}\times P_{2}\|Q_{1}\times Q_{2})=R_{2}(P_{1}\|Q_{1})\cdot R_{2}(P_{2}\|Q_{2})$ holds exactly, and the chi-squared divergence satisfies $\chi^{2}(P\|Q)=R_{2}(P\|Q)-1$.
Definition Q.11 (Chi-Squared Divergence).
The chi-squared divergence between distributions $P$ and $Q$ (with $P\ll Q$) is:
$\chi^{2}(P\|Q)=\sum_{x\in\Omega}\frac{(P(x)-Q(x))^{2}}{Q(x)}=R_{2}(P\|Q)-1$
where the last equality holds by the definition of $R_{2}$ (order-2 Rényi divergence).
@@
For $\epsilon\geq 0$, the one-sided smooth Rényi divergence is:
$R_{\alpha}^{\epsilon}(P\|Q)=\min_{P^{\prime}}R_{\alpha}(P^{\prime}\|Q),\quad P^{\prime}:\Delta(P^{\prime},P)\leq\epsilon$
where $\Delta$ denotes statistical distance. We use the one-sided variant (smoothing over $P^{\prime}$ only, $Q$ fixed), as required by Lemma Q.13.
@@
Let $\Pi_{\mathsf{IH}}$ be threshold ML-DSA with Irwin-Hall nonces (sum of $|S|$ uniforms), and $\Pi_{\mathsf{U}}$ be the hypothetical scheme with uniform nonces. Then for any PPT adversary $\mathcal{A}$ making at most $q_{s}$ signing queries:
$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}(\mathcal{A})\leq\bigl(R_{2}^{\mathsf{vec}}\bigr)^{q_{s}/2}\cdot\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}(\mathcal{A})}+q_{s}\cdot\epsilon$
where $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the full-vector per-session Rényi divergence, and the per-coordinate $R_{2}^{\epsilon,\mathsf{coord}},\epsilon$ are computed below. The $(R_{2}^{\mathsf{vec}})^{q_{s}/2}$ factor accounts for multi-session composition; see Remark Q.17. For the direct bound (Corollary Q.24), this factor is $(1+\chi^{2}_{\max})^{n\ell\,q_{s}/2}$ and the security loss grows as $\approx 6.1\times 10^{-3}\,q_{s}$ bits for $|S|=17$.
Proof Q.16.
@@
In $\Pi_{\mathsf{IH}}$ (real): $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ where each
$\hat{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$
is party $i$’s constant-term contribution in the nonce DKG.
@@
Let $m=|S|$ denote the signing set size and $B=2\lfloor\gamma_{1}/|S|\rfloor$ the per-party full nonce range (distinct from the ring degree $n=256$ used elsewhere; note $B$ here equals $2\times$ the half-range $B^{\prime}$ in Definition Q.7, i.e., $B^{\prime}=B/2=\lfloor\gamma_{1}/|S|\rfloor$). For $m$ uniforms on half-range $B/2$ with shift $\delta=\|(c\mathbf{s}_{1})_{j}\|_{\infty}\leq\beta$ (where $c$ is the challenge polynomial), specializing to $\alpha=2$ (where the factor $\alpha(\alpha-1)/2=1$):
$R_{2}^{\epsilon}\leq 1+\frac{m^{2}\delta^{2}}{B^{2}}+O(1/m!)$